OpenStudy (anonymous):

f(x)= x-x^3 find the 3 x-intercepts..

OpenStudy (anonymous):

f(x) = 0 when x = 1,0,0

OpenStudy (anonymous):

1,0,-1

hero (hero):

x-x^3 = 0 x(1-x^2) = 0 x = 0 1 - x^2 = 0 x^2 = 1 x = +/- 1 x = -1,0,1